OpenStudy (lena772):

Find the derivative of g(x)=sqrt(5-3x)

OpenStudy (shamil98):

\[\sqrt{5-3x} = (5-3x)^{\frac{ 1 }{ 2 }}\] chain rule i think

OpenStudy (shamil98):

Are you familiar with the chain rule?

OpenStudy (isaiah.feynman):

\[g(x) = \sqrt{5-3x}\] \[g'(x)= -\frac{ 3 }{ 2 }(5-3x)^{-\frac{ 3 }{ 2 }}\]

OpenStudy (shamil98):

No @Isaiah.Feynman

OpenStudy (lena772):

d dx [f(u)] = f'(u) u'

OpenStudy (shamil98):

Yes lena

OpenStudy (shamil98):

Take the derivative of (5-3x) and (5-3x)^1/2 and multiply

OpenStudy (lena772):

d (5-3x) dx

OpenStudy (lena772):

d (5-3x)^1/2 dx

OpenStudy (lena772):

These equal something?

OpenStudy (lena772):

Right?

OpenStudy (shamil98):

Yes, take the derivatives of (5-3x)^1/2 and (5-3x) then just multiply

OpenStudy (shamil98):

You result with \[g'(x) = \frac{ (5-3x)^{\frac{ -1 }{ 2 }} }{ 2 } \times (-3)\]

OpenStudy (shamil98):

Simplifying.. \[g'(x) = \frac{ -3 }{ 2(5-3x)^{1/2} } = \frac{ -3 }{ 2\sqrt{5-3x} }\]

OpenStudy (lena772):

that's all?

OpenStudy (shamil98):

Yeah.

OpenStudy (shamil98):

https://www.math.ucdavis.edu/~kouba/CalcOneDIRECTORY/chainrulesoldirectory/ChainRuleSol.html#SOLUTION 1 Take a look at this, it'll help understand the chain rule and such. It's helped me learn it in the past ^_^